# PriorDiffusion: Leverage Language Prior in Diffusion Models for Monocular Depth Estimation

### v3

@@ Abstract
Traditional monocular depth estimation suffers from inherent ambiguity and visual nuisance. We argue that language prior can enhance monocular depth estimation by leveraging the inductive bias learned during the text-to-image pre-training of diffusion models. The ability of these models to generate images that align with text indicates that they have learned the spatial relationships, size, and shape of specified objects, which can be applied to improve depth estimation. Thus, we propose PriorDiffusion, using a pre-trained text-to-image diffusion model that takes both images and corresponding text descriptions to infer affine-invariant depth through a denoising process. We also show that language prior enhances the model’s perception of specific regions of images that users care about and describe. Simultaneously, language prior acts as a constraint to accelerate the convergence of both training and the inference diffusion trajectory. By training on HyperSim and Virtual KITTI, we achieve faster training convergence, fewer inference diffusion steps, and state-of-the-art zero-shot performance across NYUv2, KITTI, ETH3D, and ScanNet. Code will be released upon acceptance.
Figure 1: Language prior in diffusion models enhances depth estimation by leveraging the geometric prior aligned with the language description. It is learned during text-to-image generation pre-training, as to generate images under different viewpoints and layouts that accurately reflect the text, the model must comprehend the size and shape of specified objects, their spatial relationship, and the scale of the scene.
@@ 1 Introduction
Monocular depth estimation requires the model to predict pixel-wise depth from a single image. Recent advancements in diffusion-based models have shown promise for generating high-quality depth maps [11, 29, 57, 58, 83, 30, 17, 24, 15, 21, 25, 65] due to their ability to capture complex structures and fine details by progressively denoising a latent representation to produce accurate depth predictions. However, it remains a challenging task due to inherent ambiguities and various visual nuisances. Unlike stereo or multi-view depth estimation methods, monocular approaches lack depth cues provided by multiple perspectives, leading to difficulty in accurately discerning the relative sizes, distances, and spatial relationships of objects within a scene. That is, different real-world 3D scenes can project to the same 2D image, causing inherent ambiguities. For example, texture ambiguity. Surfaces that are uniformly textured or have repetitive patterns, like tiled floors, can create confusion for depth estimators. The models might misinterpret the repetition as uniform depth, even though the actual depth might vary (e.g., a tiled floor receding into the distance). Also, visibility ambiguity creates limitations for discerning small, partial or occluded objects purely from the visual signal. On the other hand, visual nuisances such as illumination, motion blur, atmospheric conditions, and perspective distortion distort or obscure critical visual cues, leading vision algorithms to misinterpret scene structure, depth, and object boundaries, thus reducing their accuracy and reliability.
@@
Shown in Figure 1, an emerging solution lies in leveraging language as a complementary modality to provide a prior to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of priors about spatial relationships, scene types, object sizes, and depth hierarchies. In cases where an object is small, partially occluded, visually similar to the background, or its type is ambiguous from a single image, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling”, the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities. As shown in Figure 2, better perception can be achieved for insignificant regions indicated in the language description.
@@
To this end, we propose PriorDiffusion, incorporating human-provided language prior of the scene as an inductive bias to guide the depth map prediction. During text-to-image pre-training, diffusion models learn to generate diverse images under various viewpoints and scene layouts that align with the provided language descriptions. The ability of these models to generate images that align with text suggests they have learned the spatial relationships, size, and shape of specified objects in the described 3D scene, which can be leveraged to improve depth estimation. In our PriorDiffusion, during the denoising process, both the image and the text input are used by the model as conditions to predict the noise to be removed. Ultimately, Gaussian noise is progressively refined into a depth map that aligns with both the input image and the language description. In real-world applications when embodied agents are integrated with PriorDiffusion, humans can describe the environment to help agents better perceive it. When generalizing to a new environment, language, having less covariate shift compared to images, allows humans to describe the new environment, aiding agents in better adapting and generalizing to it.
To support our hypothesis, we conduct training on two synthetic datasets, HyperSim [53] and Virtual KITTI [3], and conduct zero-shot evaluation on four real-world datasets, NYUv2 [60], KITTI [22], ETH3D [59], and ScanNet [9]. Given the difficulty of obtaining sufficient human-provided text descriptions for every training and inference image, we utilize a vision-language model, i.e. LLaVA [10], to generate descriptions for each image, simulating human annotation. we achieve faster training convergence, fewer inference diffusion steps, and state-of-the-art zero-shot performance across NYUv2, KITTI, ETH3D, and ScanNet.
Our contributions. By leveraging language prior for diffusion-based monocular depth estimation, we demonstrate that:
@@
  • Language prior provides an inductive bias about the existence, geometric properties, and spatial relationships of objects specified in the language description, allowing depth estimators to better perceive the depth of those objects.
@@
  • Language prior directs the model’s attention to specific regions, guiding it to perceive the 3D scene according to the user’s intent. This is particularly beneficial for regions that pose challenges to vision systems, such as those with small size, poor illumination, occlusion, or high visual similarity to the background.
@@
  • Language prior serves as a constraint to speed up training convergence, and serves as a good initialization of diffusion trajectory to speed up inference. This improves the efficiency of the diffusion model, optimizing both its performance and computational resource usage.
Figure 2: 3D reconstruction. For insignificant regions indicated in language description, PriorDiffusion achieves better depth estimates and reconstructions thanks to the use of language.
@@
Figure 3: Pipeline of PriorDiffusion. We train the diffusion model to predict the noise added into the noisy depth latent $\mathbf{z}_{t}$ at the time step $t$, based on $\mathbf{z}_{t}$, the input image $x$, and the corresponding language description $c$. During inference, the diffusion model predicts noise for $\mathbf{z}_{t}$ at each time step and gradually denoise it from $\mathbf{z}_{T}$ (pure Gaussian noise) into $\mathbf{z}_{0}$ (pure depth latent). Then $\mathbf{z}_{0}$ is decoded into the depth prediction using a frozen variational decoder.
@@ 2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [2, 16, 32, 39, 64, 73, 81, 74] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [76, 68, 51, 50, 30, 69, 72, 12, 71, 38], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [76] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [68] learns from large-scale automatically annotated data. DPT [51] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [50] mixes multiple datasets with training objectives invariant to depth range and scale. Depth Anything [68] and Depth Anything v2 [69] train a depth foundation model using large-scale unlabeled data. Existing methods try to predict depth solely from images, which suffers from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. On the other hand, our PriorDiffusion leverages language prior in Diffusion Models to provide a thorough understanding of the scene associated with the input text, to resolve the ambiguity and the nuisance variables in the visual algorithm.
@@
Diffusion model for monocular depth estimation. Denoising Diffusion Probabilistic Model (DDPM) [26] learns to reverse a diffusion process that progressively degrades images with Gaussian noise so that they can draw samples from the data distribution by applying the reverse process to random noise. They have been applied to various tasks like text-to-image generation [54, 49, 89, 88, 63], super resolution [55, 35, 20], image inpainting [42, 8, 70], 3D object generation [46, 34, 41], etc. Several approaches have explored the use of DDPM for metric depth estimation. DiffusionDepth [11] learns to denoise random depth distribution into a depth map with monocular visual conditions. Similarly, DDP [29] encodes the input image and use diffusion to decodes it into a depth map. DepthGen [57] extends a multi-task diffusion model for metric depth prediction. Its successor, DDVM [58], focuses on pretraining with synthetic and real datasets to improve depth estimation performance. Build upon the diffusion model, DepthFM [24] uses flow matching to conduct a faster monocular depth estimation. In light of strong representation learned by pre-trained, text-to-image diffusion models, recent methods like VPD [83], E2E FT [21], GeoWizard [17], Lotus [25], GenPercept [65], and Marigold [30] utilize a pretrained Stable Diffusion model [54] as a backbone to predict depth through a denoising process. Specifically, Marigold [30] concatenates input images with diffusion latent as conditions for denoising U-Net to predict noise that needs to be removed, to gradually denoise a Gaussian noise to a depth map. While previous works only focused on using images as conditions of diffusion models to predict depth, our PriorDiffusion utilizes language prior as inductive bias for diffusion models to resolve ambiguity and visual nuisances, and conduct 3D perception in a controllable way.
Figure 4: Visualization on NYUv2. Compared to Margold, our PriorDiffusion demonstrates more accurate depth prediction for a given input image, particularly for instances specified in the language description (marked in red). This is achieved by leveraging the language prior within the text-to-image diffusion model, enhancing the model’s ability to comprehend the geometric characteristics of these objects. Additionally, the language description directs the model’s attention to relevant regions, especially the ones that are easily neglected by visual signals, like “a soap dispenser” in the first row. It enables the depth estimator to perceive the 3D scene in a more controlled manner that aligns with the user’s interest.
@@
Language for monocular depth estimation. Vision-Language models  [4, 37, 36, 45, 48] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [66, 67, 79, 6, 80, 81, 87, 74, 75, 43, 62, 31]. For example, CLIP [48] use contrastive learning on text-image pairs, which allows for various applications. These include few-shot image classification  [19, 78, 47], object detection  [52, 86], image segmentation  [52, 84], and 3D perception [28, 80, 81, 87, 85, 77, 7]. Given their growing capabilities, some research [1, 28, 81, 82] has focused on applying vision-language models to monocular depth estimation. Specifically, DepthCLIP [81] utilizes CLIP’s semantic depth response as an ordinal relationship alongside a depth projection method to conduct zero-shot monocular depth estimation. WorDepth [74] learns the variational prior distribution of 3D scenes from text. RSA [75] predicts scale using language description to align relative depth to metric scale. While language has been applied in various aspects of 3D perception, our work is the first to use language prior for diffusion models to enhance 3D perception.
3 Method

### v4

# Iris: Integrating Language into Diffusion-based Monocular Depth Estimation

@@ Abstract
Traditional monocular depth estimation suffers from inherent ambiguity and visual nuisances. We demonstrate that language can enhance monocular depth estimation by providing an additional condition (rather than images alone) aligned with plausible 3D scenes, thereby reducing the solution space for depth estimation. This conditional distribution is learned during the text-to-image pre-training of diffusion models. To generate images under various viewpoints and layouts that precisely reflect textual descriptions, the model implicitly models object sizes, shapes, and scales, their spatial relationships, and the overall scene structure. In this paper, Iris, we investigate the benefits of our strategy to integrate text descriptions into training and inference of diffusion-based depth estimation models. We experiment with three different diffusion-based monocular depth estimators (Marigold, Lotus, and E2E-FT) and their variants. By training on HyperSim and Virtual KITTI, and evaluating on NYUv2, KITTI, ETH3D, ScanNet, and DIODE, we find that our strategy improves the overall monocular depth estimation accuracy, especially in small areas. It also improves the model’s depth perception of specific regions described in the text. We find that by providing more details in the text, the depth prediction can be iteratively refined. Simultaneously, we find that language can act as a constraint to accelerate the convergence of both training and the inference diffusion trajectory. Code and generated text data will be released upon acceptance.
footnotetext: * Equal contribution.
Figure 1: Integrating language into diffusion models enhances monocular depth estimation by providing an additional condition (rather than images alone) associated with plausible 3D scenes, thus reducing the solution space for depth estimation. This conditional distribution is initially learned during text-to-image generation pre-training of diffusion models, as to generate images under different viewpoints and layouts that accurately reflect the text, the model needs to implicitly model the size and shape of specified objects, their spatial relationship, and the structure of the scene. Then the conditional distribution is associated with plausible 3D scenes during fine-tuning with image-text-depth pairs.
@@ 1 Introduction
Monocular depth estimation requires the model to predict pixel-wise depth from a single image. Recent advancements in diffusion-based models have shown promise for generating high-quality depth maps [duan2023diffusiondepth, ji2023ddp, saxena2023monocular, saxena2024surprising, zhao2023unleashing, Marigold, fu2024geowizard, gui2024depthfm, fei2019geo, garcia2024fine, he2024lotus, xu2024diffusion] due to their ability to capture complex structures and fine details by progressively denoising a latent representation to produce accurate depth predictions. However, it remains a challenging task due to inherent ambiguities and various visual nuisances. Unlike stereo or multi-view depth estimation methods, monocular approaches lack visual features provided by multiple perspectives, leading to difficulty in accurately discerning the relative sizes, distances, and spatial relationships of objects within a scene. That is, different real-world 3D scenes can project to the same 2D image, causing inherent ambiguities. For example, texture ambiguity. Surfaces that are uniformly textured or have repetitive patterns, like tiled floors, can create confusion for depth estimators. The models might misinterpret the repetition as uniform depth, even though the actual depth might vary (e.g., a tiled floor receding into the distance). Also, visibility ambiguity creates limitations for discerning small, partial or occluded objects purely from the visual signal. On the other hand, visual nuisances such as illumination, motion blur, atmospheric conditions, and perspective distortion distort or obscure critical visual cues, leading vision algorithms to misinterpret scene structure, depth, and object boundaries, thus reducing their accuracy and reliability.
Figure 2: Language improves the depth perception of specified insignificant (and potentially ambiguous) regions.
@@
Shown in Figure 1, an emerging solution lies in integrating language as a complementary modality to provide an additional condition to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of conditions about spatial relationships, scene types, object sizes, and depth hierarchies. In cases where an object is small, partially occluded, visually similar to the background, or its type is ambiguous from a single image, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling,” the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities. As shown in Figure 2, better depth perception can be achieved for regions specified in the language description. Without language, the baseline model misinterprets the background kitchen as a flat surface, likely a wallpaper or a painting. With the additional condition “a kitchen in the background,” the model applies an additional condition about the structure of a kitchen, leading to more accurate depth prediction. In real-world applications of embodied agents, human provides text instruction to agents, which contains the description of the scene and can be used to enhance models’ 3D perception at no additional cost. In the example of Figure 2, if an agent is instructed to grab something “in the kitchen in the background”, the instruction can be integrated into the 3D perception model to avoid treating the background as a flat surface. Also, when generalizing to a new environment, language, having less covariate shift compared to images [zeng2024rsa, zeng2024wordepth], allows humans to describe the new environment, aiding agents in better adapting and generalizing to it.
@@
To this end, we investigate incorporating language descriptions of the scene as conditions to enhance the depth map prediction. During text-to-image pre-training, diffusion models learn to generate diverse images under various viewpoints and scene layouts that align with the provided language descriptions. The ability of these models to generate images that align with text suggests these models are implicitly modeling the spatial relationships, size, shape, and scales of specified objects, as well as the structure of the 3D scene, then associate such implicit modeling with text. When integrating text into the training and inference of diffusion-based monocular depth estimator, text is used as a condition to reduce the solution space of possible 3D scenes, thus improve the accuracy of depth prediction. In our implementation, during the denoising process, both the image and the text input are used by the model as conditions to predict the noise to be removed. Ultimately, Gaussian noise is progressively refined into a depth map that aligns with both the input image and the language description.
To support our hypothesis, we examine three different diffusion-based monocular depth estimator and their variants: Marigold [Marigold], Lotus [he2024lotus], and E2E-FT [garcia2024fine]. We conduct training on two synthetic datasets, HyperSim [roberts2021hypersim] and Virtual KITTI [VirtualKITTI], and conduct zero-shot evaluation on five real-world datasets, NYUv2 [NYU-Depth-V2], KITTI [KITTI], ETH3D [ETH3D], ScanNet [ScanNet], and DIODE [DIODE]. Given the difficulty of obtaining sufficient human-provided text descriptions for every training and inference image, we utilize a vision-language model, i.e. LLaVA v1.6 [liu2023improved] and InternVL3-8B [zhu2025internvl3], to generate descriptions for each image, simulating human annotation. By integrating text into both the training and inference stages of a diffusion-based depth estimator, we find that it enhances overall monocular depth estimation accuracy and improves the model’s depth perception in regions described by the text. Furthermore, providing more detailed textual descriptions enables iterative refinement of the predicted depth. At the same time, language serves as a constraint that accelerates the convergence of both training and the diffusion trajectory during inference. Interestingly, we also observe that in some cases, even when text is only provided in training or inference time, the accuracy of monocular depth estimation still shows a slight improvement. One possible explanation is that when the model is trained with text but evaluated without it, the training-time text descriptions help the model disambiguate challenging or visually ambiguous regions. By learning these associations, the model develops a stronger correspondence between such regions and their ground-truth depth, which in turn improves depth prediction even when text is absent at inference time. Conversely, when the model is trained without text but evaluated with it, the diffusion model can still partially utilize text features. Although no text–depth correspondence is learned during training, text prompts can activate conditions that diffusion models have learned from large-scale text-to-image pre-training—such as associations with appearance, structure, or texture. These implicit priors can provide weak but useful guidance, thereby improving depth estimation to some extent even without explicit text–depth supervision.
Our findings. We summarize our findings of the influence of integrating language into the diffusion-based monocular depth estimator as below:
@@
  • Finding 1: Language provides a condition about the existence, geometric properties, and spatial relationships of objects and scene structures, helping depth estimators to reduce the depth solution space to better perceive depth, especially in insignificant or ambiguous regions.
@@
  • Finding 2: Depth prediction can be iteratively refined with more details in the text description. This is particularly beneficial for regions that pose challenges to vision systems, such as those with small size, poor illumination, occlusion, or high visual similarity to the background.
@@
  • Finding 3: Language serves as a constraint that accelerates training convergence and provides a good initialization of the diffusion trajectory to speed up inference. This improves the efficiency of the diffusion model.
@@
Figure 3: Pipeline to integrate text. We train the diffusion model to predict the noise added into the noisy depth latent $\mathbf{z}_{t}$ at the time step $t$, based on $\mathbf{z}_{t}$, the input image $x$, and the corresponding language description $c$. During inference, the diffusion model predicts noise for $\mathbf{z}_{t}$ at each time step and gradually denoise it from $\mathbf{z}_{T}$ (pure Gaussian noise) into $\mathbf{z}_{0}$ (pure depth latent). Then $\mathbf{z}_{0}$ is decoded into the depth prediction using a frozen variational decoder.
@@ 2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [bhat2021adabins, dorn, lao2024sub, Va-depthnet, wong2020targeted, yuan2022neural, DepthCLIP, zeng2024wordepth, chung2025eta, rim2025protodepth] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [zhang2022hierarchical, DepthAnything, DPT, midas, Marigold, DepthAnythingV2, yin2021learning, eftekhar2021omnidata, yin2020diversedepth, li2018megadepth], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [zhang2022hierarchical] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [DepthAnything] learns from large-scale automatically annotated data. DPT [DPT] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [midas] mixes multiple datasets with training objectives invariant to depth range and scale. Depth Anything [DepthAnything] and Depth Anything v2 [DepthAnythingV2] train a depth foundation model using large-scale unlabeled data. Existing methods try to predict depth solely from images, which suffer from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. In this paper, we investigate integrating language into diffusion-based monocular depth estimators to provide conditions of 3D scenes, to resolve the ambiguity and the nuisance variables in the pure visual algorithm.
@@
Diffusion model for monocular depth estimation. Denoising Diffusion Probabilistic Model (DDPM) [DDPM] learns to reverse a diffusion process that progressively degrades images with Gaussian noise so that they can draw samples from the data distribution by applying the reverse process to random noise. They have been applied to various tasks like text-to-image generation [StableDiffusion, ramesh2022hierarchical, zhu2024boundary, zhu2022discrete, wang2023diffusion], super resolution [saharia2022image, li2022srdiff, gao2023implicit], image inpainting [lugmayr2022repaint, corneanu2024latentpaint, yang2023uni], 3D object generation [qian2023magic123, li20223ddesigner, liu2024one], etc. Several approaches have explored the use of DDPM for metric depth estimation. DiffusionDepth [duan2023diffusiondepth] learns to denoise random depth distribution into a depth map with monocular visual conditions. Similarly, DDP [ji2023ddp] encodes the input image and use diffusion to decodes it into a depth map. DepthGen [saxena2023monocular] extends a multi-task diffusion model for metric depth prediction. Its successor, DDVM [saxena2024surprising], focuses on pretraining with synthetic and real datasets to improve depth estimation performance. Build upon the diffusion model, DepthFM [gui2024depthfm] uses flow matching to conduct a faster monocular depth estimation. In light of strong representation learned by pre-trained, text-to-image diffusion models, recent methods like VPD [zhao2023unleashing], E2E-FT [garcia2024fine], GeoWizard [fu2024geowizard], Lotus [he2024lotus], GenPercept [xu2024diffusion], Jasmine [wang2025jasmine], and Marigold [Marigold] utilize a pretrained Stable Diffusion model [StableDiffusion] as the backbone to predict depth through a denoising process. Specifically, Marigold [Marigold] concatenates input images with diffusion latents as conditions for the denoising U-Net to predict the noise to be removed, gradually denoising Gaussian noise into a depth map. While previous works have primarily focused on using images as the conditioning input to diffusion models for depth prediction, we investigate use language as an an additional condition to improve monocular depth estimation.
@@
Language for monocular depth estimation. Vision-Language models  [Dino, Blip-2, Blip, Dinov2, CLIP] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [yang2024unitouch, yang2024neurobind, DSPoint, iQuery, PointCLIP, DepthCLIP, PointCLIP_v2, zeng2024wordepth, zeng2024rsa, ni2025efficient, tong2024metamorph, kim2024openvla, ni2025homer]. For example, CLIP [CLIP] use contrastive learning on text-image pairs, which allows for various applications. These include few-shot image classification  [gao2021clip_cls, zhang2021tip_cls, VT-CLIP], object detection  [rao2021denseclip_detect_seg, Detic], image segmentation  [rao2021denseclip_detect_seg, zhou2021denseclip_seg], and 3D perception [DepthCLIPv2, PointCLIP, DepthCLIP, PointCLIP_v2, zhou2023uni3d, zhang2023clip, chen2023clip2scene]. Given their growing capabilities, some research [DepthCLIP_auty, DepthCLIPv2, DepthCLIP, VPD] has focused on applying vision-language models to monocular depth estimation. Specifically, DepthCLIP [DepthCLIP] utilizes CLIP’s semantic depth response as an ordinal relationship alongside a depth projection method to conduct zero-shot monocular depth estimation. WorDepth [zeng2024wordepth] learns the variational prior distribution of 3D scenes from text. RSA [zeng2024rsa] predicts scale using language description to align relative depth to metric scale. Although language has been incorporated into several areas of 3D perception, its influence on diffusion-based monocular depth estimation has yet to be thoroughly investigated.
3 Formulation